How many moles of KCl are in 250.0 mL of 0.028 M KCl?
1
Expert's answer
2020-11-20T10:08:29-0500

number of moles = ?

volume of KCl = 250.0 mL = 250.0/1000 L = 0.25 L

molar concentration of KCl = 0.028 M


number of moles = volume × molar concentration

number of moles = 0.25 × 0.028 = 0.007 mol.